# Score-based Neural Ordinary Differential Equations for Computing Mean Field Control Problems

### v1

@@ 4.2 Modified HJB equation for MFC
In this section, we present a system of two PDEs that characterize the optimal solution for the MFC problem. Different from the traditional FP-HJB pair (see (Bensoussan et al., 2013, Chapter 4)), our system consists of a transport equation obtained from equation 24, and a modified HJB equation, tailored for the modified MFC problem equation 23. This characterization could serve as a regularizer to enhance the loss function in numerical algorithms. We define the Hamiltonian $H:\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ by
@@
This definition aligns with classical control theory and is closely related to the maximum principle (Zhou & Lu, 2023). The solution of the MFC problem is summarized by the following proposition.
@@
Let $L$ be strongly convex in $v$, then the solution to the MFC problem equation 23 is as follows. Consider a function $\psi\colon[0,t_{\textrm{end}}]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, such that
@@
$\inf_{f}\int_{0}^{t_{\text{end}}}\int_{\mathbb{R}^{d}}\frac{1}{2}\left|f(t,x)+\gamma\nabla_{x}\log\rho(t,x)\right|^{2}\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t+\int_{\mathbb{R}^{d}}G(x,\rho(t_{\text{end}},x))\rho(t_{\text{end}},x)\,\mathrm{d}x\,,$ (28)
subject to equation 24. In this case, the result becomes the Corollary below.
@@
In the LQ problem where $L(t,x,v)=\frac{1}{2}|v|^{2}$ and $F(t,x,\rho)=0$, the solution of the MFC problem equation 28 is given by
@@
If we further define the Fisher information as $I[\rho]:=\int_{\mathbb{R}^{d}}\left|\nabla_{x}\log\rho(x)\right|^{2}\rho(x)\,\mathrm{d}x$, then the equation for $\psi$ in equation equation 29 becomes
@@
Let $z_{t}$ be the characteristic line of the state trajectory, satisfying $\partial_{t}z_{t}=f(t,z_{t})$ where $f$ is given by equation 26. Then, a residual of the HJB equation along the trajectory $z_{t}$ can be computed to enhance the objective function. For example, in the LQ example in Corollary 1, we know that $\psi$ satisfies
@@
Also, by equation 22, we can recover the optimal velocity field $v$ through
@@
where $f$ is the optimal vector field of the modified MFC problem. This regularization technique could be extended to a more general setting, such as the flow matching problem for overdamped Langevin dynamics. See Corollary 2 in Appendix A for details.
@@ 4.3 Numerical algorithms
In this section, we present numerical algorithms to solve the MFC problem. We minimize the objective equation 23 to obtain the optimal vector field $f$. Additionally, we can incorporate the residual of the HJB equation as a regularizer to enhance the loss functional, as discussed after Corollary 1.
The composed velocity field $f$ is parametrized as a neural network, as defined in equation 8. In order to simulate the cost functional numerically, we sample multiple initial points $z_{0}^{(n)}\sim\rho_{0}$ with a batch size $N_{z}$. Then, we can simulate the dynamics of $z_{t}^{(n)}$, $l_{t}^{(n)}$, $\tilde{l}_{t}^{(n)}$, $s_{t}^{(n)}$, and $H_{t}^{(n)}$ numerically for each particle through equation 9, equation 10, equation 11, equation 12, and equation 13 respectively, where the derivatives of $f$ are obtained via auto-differentiation. Then, we simulate the cost functional equation 23 by
$\mathcal{L}_{\text{cost}}=\dfrac{1}{N_{z}}\sum_{n=1}^{N_{z}}\left[\sum_{j=0}^{N_{t}-1}\left(L\left(t_{j},z_{t_{j}}^{(n)},f(t_{j},z_{t_{j}}^{(n)};\theta)-\gamma s_{t_{j}}^{(n)}\right)+F\left(t_{j},z_{t_{j}}^{(n)},\tilde{l}_{t_{j}}^{(n)}\right)\right)\Delta t+G(z_{t_{N_{t}}}^{(n)},\tilde{l}_{t_{N_{t}}}^{(n)})\right]\,.$ (31)
@@
MFC problem equation 19 equation 20, $N_{t}$, $N_{z}$, network structure equation 8, learning rate, number of iterations
@@
         compute $\left(\nabla_{z},\nabla_{z}\cdot,\nabla_{z}(\nabla_{z}\cdot)\right)f(t_{j},z_{t_{j}}^{(n)};\theta)$ according to equation 8
         compute $z_{t_{j+1}}^{(n)}$, $\tilde{l}_{t_{j+1}}^{(n)}$, $s_{t_{j+1}}^{(n)}$ through the forward Euler scheme equation 9, equation 11, and equation 12
@@
To further improve performance, we can add a penalty of the residual of the HJB equation, denoted by $\mathcal{L}_{\text{HJB}}$, to enhance the loss function. This regularization technique is detailed in Appendix B.1. The total loss function is hence given by
@@ 5 Numerical results
We present numerical results for several examples in this section, including the regularized Wasserstein proximal operator (RWPO), the flow matching problem, a linear-quadratic (LQ) problem with an entropy potential cost, and an example with double well potential. For examples with exact solutions, we present all errors for the density ($\text{err}_{\rho}$), velocity field ($\text{err}_{f}$), and score function ($\text{err}_{s}$). These errors are calculated as averages over $10$ independent runs, see equation 95 in Appendix B.3 for details. The parameters used for all numerical examples are provided in Appendix B.4. Part of the results are deferred to Appendix C.
@@ 5.1 Regularized Wasserstein proximal operator
In the RWPO problem, the objective is a regularized Benamou-Brenier formulation for optimal transportation (Benamou & Brenier, 2000b). We minimize the cost functional
@@
subject to the FP equation $\partial_{t}\rho(t,x)+\nabla_{x}\cdot(\rho(t,x)v(t,x))=\gamma\Delta_{x}\rho(t,x)$. Here we set $G(x)=|x|^{2}/2$ and $\rho_{0}(x)=(8\pi)^{-d/2}\exp\left(-|x|^{2}/8\right)$. We test Algorithm 1 on this problem in $1$, $2$, and $10$ dimensions. These errors are summarized in Table 2. Detailed definitions for $\text{err}_{\rho}$, $\text{err}_{f}$, and $\text{err}_{s}$ are given in Appendix B.3. Additionally, we report the cost gap, which represents the difference between the computed cost and the optimal cost, averaged over 10 independent runs.
@@
The results are also visualized in Figure 1. The plot on the left shows the cost functional through training in $1$ dimension, which becomes close to the optimal cost in red. The plot in the middle compares the evolution of the density function computed through equation 11 under trained velocity with the true density evolution. Our density dynamic equation 4 accurately captures the density evolution. The plot on the right shows the particle trajectories of $z_{t}$ in $2$ dimensions and compares them with the stochastic dynamics. Our probability flow ODE demonstrates a structured behavior. We leave further numerical results of this example in Appendix C.1.
Figure 1: Numerical results for the RWPO problem. Left: cost functional through training with optimal cost in $1d$. Middle: density evolution through equation 4 and comparison with true density in $1d$. Right: trajectories of $z_{t}$ and comparison with the stochastic dynamic. The score dynamic demonstrates a structured behavior compared with the corresponding stochastic trajectory simulated by stochastic differential equations.
@@ 5.2 Flow matching for solving FP equations
Flow matching has emerged as an important problem, which is closely related to generative models (Lipman et al., 2022; Boffi & Vanden-Eijnden, 2023). The problem is to simulate the probability density function of a stochastic dynamics:
@@
where $b\colon\mathbb{R}^{+}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a known drift vector field, $\rho_{0}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is an initial value probability density function. In this example, we also assume the drift function satisfies $b(t,x)=-\nabla V(x)$, which is the negative gradient of some potential function $V(x)\in C_{\text{loc}}^{4}(\mathbb{R}^{d})$. In this case, SDE equation 33 is the overdamped Langevin dynamic. To simulate the density function of stochastic process $X_{t}$ in equation 33, we design the following MFC problem. We minimize the objective functional (loss function) as
@@
subject to the FP equation equation 20. After the score transformation equation 22, the above second order MFC problem becomes
@@
subject to the transport equation equation 24. The state dynamic hence becomes $\partial_{t}z_{t}=-\nabla_{z}V(z_{t})-\gamma s_{t}$. This transformation equation 22 seems to complicate the problem. However, we are able to learn the whole FP equation of the overdamped Langevin dynamic with a given initial distribution $\rho_{0}$ or samples $\{z_{0}^{(n)}\}_{n=1}^{n_{z}}$. More importantly, we can record all intermediate time steps and compute the density function during the entire time domain $[0,t_{\text{end}}]$.
We first present the flow matching problem for an Ornstein–-Uhlenbeck (OU) process, where an explicit solution is available, given in Appendix A.3. The algorithm is described with details in Appendix B.1 and summarized in Algorithm 3. Similar to the previous section, we compare the numerical results with and without regularizations, presented in Table 4. The errors for the regularized algorithm are significantly smaller. We also provide visualized results in Figure 3 and 4 in Appendix C.2, which demonstrates the accuracy of our algorithm.
@@
In addition, we design an algorithm that partitions a long time interval into several sub-intervals, which could potentially resolve the issue of long time horizon. The model is trained consecutively over each sub-interval, referred to as the multi-stage splicing method. We present an example in $2$ dimensions to demonstrate effectiveness of this splicing method, in which the invariant distribution of the SDE equation 33 is in a double moon shape (cf. equation 99). Our splicing method (Algorithm 4) is able to capture the state dynamic within each interval, and pass the information to the next interval. The detailed numerical implementation and numerical results are presented in Appendix C.2.
We also present two more examples in Appendix C.3 and C.4, where we apply the proposed neural ODE system to approximate the LQ MFC problems.

### v2

@@ 4.2 Modified HJB equation for MFC
In this section, we present a system of two PDEs that characterize the optimal solution for the MFC problem. Different from the traditional FP-HJB pair (see [13, Chapter 4]), our system consists of a transport equation obtained from (24), and a modified HJB equation, tailored for the modified MFC problem (23). This characterization could serve as a regularizer to enhance the loss function in numerical algorithms. We define the Hamiltonian $H:\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ by
@@
This definition aligns with classical control theory and is closely related to the maximum principle [32]. The solution of the MFC problem is summarized by the following proposition.
@@
Let $L$ be strongly convex in $v$, then the solution to the MFC problem (23) is as follows. Consider a function $\psi\colon[0,t_{\textrm{end}}]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, such that
@@
$\inf_{f}\int_{0}^{t_{\text{end}}}$ $\int_{\mathbb{R}^{d}}\frac{1}{2}\left|f(t,x)+\gamma\nabla_{x}\log\rho(t,x)\right|^{2}\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t$ (28)
subject to (24). In this case, the result becomes the Corollary below.
@@
In the LQ problem where $L(t,x,v)=\frac{1}{2}|v|^{2}$ and $F(t,x,\rho)=0$, the solution of the MFC problem (28) is given by
@@
If we further define the Fisher information as $I[\rho]:=\int_{\mathbb{R}^{d}}\left|\nabla_{x}\log\rho(x)\right|^{2}\rho(x)\,\mathrm{d}x$, then the equation for $\psi$ in (29) becomes
@@
Let $z_{t}$ be the characteristic line of the state trajectory, satisfying $\partial_{t}z_{t}=f(t,z_{t})$ where $f$ is given by (26). Then, a residual of the HJB equation along the trajectory $z_{t}$ can be computed to enhance the objective function. For example, in the LQ example in Corollary 1, we know that $\psi$ satisfies
@@
Also, by (22), we can recover the optimal velocity field $v$ through
@@
where $f$ is the optimal vector field of the modified MFC problem. This regularization technique could be extended to a more general setting, such as the flow matching problem for overdamped Langevin dynamics. See Corollary 2 in A for details.
@@ 4.3 Numerical algorithms
In this section, we present numerical algorithms to solve the MFC problem. We minimize the objective (23) to obtain the optimal vector field $f$. Additionally, we can incorporate the residual of the HJB equation as a regularizer to enhance the loss functional, as discussed after Corollary 1.
The composed velocity field $f$ is parametrized as a neural network, as defined in (8). In order to simulate the cost functional numerically, we sample multiple initial points $z_{0}^{(n)}\sim\rho_{0}$ with a batch size $N_{z}$. Then, we can simulate the dynamics of $z_{t}^{(n)}$, $l_{t}^{(n)}$, $\tilde{l}_{t}^{(n)}$, $s_{t}^{(n)}$, and $H_{t}^{(n)}$ numerically for each particle through (9), (10), (11), (12), and (13) respectively, where the derivatives of $f$ are obtained via auto-differentiation. Then, we simulate the cost functional (23) by
$\mathcal{L}_{\text{cost}}=\dfrac{1}{N_{z}}\sum_{n=1}^{N_{z}}\sum_{j=0}^{N_{t}-1}$ $\left[\left(L\left(t_{j},z_{t_{j}}^{(n)},f(t_{j},z_{t_{j}}^{(n)};\theta)-\gamma s_{t_{j}}^{(n)}\right)\right.\right.$ (31)
$\left.\left.+F\left(t_{j},z_{t_{j}}^{(n)},\tilde{l}_{t_{j}}^{(n)}\right)\right)\Delta t+G(z_{t_{N_{t}}}^{(n)},\tilde{l}_{t_{N_{t}}}^{(n)})\right]\,.$
@@
MFC problem (19) (20), $N_{t}$, $N_{z}$, network structure (8), learning rate, number of iterations
@@
         compute $\left(\nabla_{z},\nabla_{z}\cdot,\nabla_{z}(\nabla_{z}\cdot)\right)f(t_{j},z_{t_{j}}^{(n)};\theta)$ according to (8)
         compute $z_{t_{j+1}}^{(n)}$, $\tilde{l}_{t_{j+1}}^{(n)}$, $s_{t_{j+1}}^{(n)}$ through the forward Euler scheme (9), (11), and (12)
@@
To further improve performance, we can add a penalty of the residual of the HJB equation, denoted by $\mathcal{L}_{\text{HJB}}$, to enhance the loss function. This regularization technique is detailed in B.1. The total loss function is hence given by
@@ 5 Numerical results
We present numerical results for several examples in this section, including the regularized Wasserstein proximal operator (RWPO), the flow matching problem, a linear-quadratic (LQ) problem with an entropy potential cost, and an example with double well potential. For examples with exact solutions, we present all errors for the density ($\text{err}_{\rho}$), velocity field ($\text{err}_{f}$), and score function ($\text{err}_{s}$). These errors are calculated as averages over $10$ independent runs, see (101) in B.3 for details. The parameters used for all numerical examples are provided in B.4.
@@ 5.1 Regularized Wasserstein proximal operator
In the RWPO problem, the objective is a regularized Benamou-Brenier formulation for optimal transportation [21]. We minimize the cost functional
@@
subject to the FP equation $\partial_{t}\rho(t,x)+\nabla_{x}\cdot(\rho(t,x)v(t,x))=\gamma\Delta_{x}\rho(t,x)$. Here we set $G(x)=|x|^{2}/2$ and $\rho_{0}(x)=(8\pi\gamma)^{-d/2}\exp\left(-|x|^{2}/(8\gamma)\right)$.
After the transformation to probability flow through (22), the problem becomes
subject to the transport equation
$\partial_{t}\rho(t,x)+\nabla_{x}\cdot(\rho(t,x)f(t,x))=0\,.$
The optimal density evolution is given by
and the optimal velocity field is
$f(t,x)=-\dfrac{x}{2(2-t)}\,.$
We test Algorithm 1 on this problem in $1$, $2$, and $10$ dimensions. These errors are summarized in Table 2. Detailed definitions for $\text{err}_{\rho}$, $\text{err}_{f}$, and $\text{err}_{s}$ are given in B.3. Additionally, we report the cost gap, which represents the difference between the computed cost and the optimal cost, averaged over 10 independent runs.
@@
The results are also visualized in Figure 1. The plot on the left shows the cost functional through training in $1$ dimension, which becomes close to the optimal cost in red. The plot in the middle compares the evolution of the density function computed through (11) under trained velocity with the true density evolution. Our density dynamic (4) accurately captures the density evolution. The plot on the right shows the particle trajectories of $z_{t}$ in $2$ dimensions and compares them with the stochastic dynamics. Our probability flow ODE demonstrates a structured behavior.
Figure 1: Numerical results for the RWPO problem. Left: cost functional through training with optimal cost in $1d$. Middle: density evolution through (4) and comparison with true density in $1d$. Right: trajectories of $z_{t}$ and comparison with the stochastic dynamic in $2d$. The score dynamic demonstrates a structured behavior compared with the corresponding stochastic trajectory simulated by stochastic differential equations.
We present additional numerical results in Figure 2. The first, second, and third rows of Figure 2 show numerical results in $1$, $2$, and $10$ dimensions, respectively. The first column shows the curves of the cost function through training. The light blue shadows plot the standard deviation observed during $10$ independent test runs. The red dashed lines represent the cost under the optimal control field. We observe that our algorithm nearly reaches the optimal cost.
The fourth column in the first row of Figure 2 plots the score function (in blue) at $t=1$ using the data points $\{(z^{(n)}_{t_{N_{t}}},s^{(n)}_{t_{N_{t}}})\}_{n=1}^{N_{z}}$. This curve coincides with the true score function in orange. Note that the score function at $t=0$ is given, so we only plot the score at a terminal time. The second and third figures in the fourth column of Figure 2 show a density plot of the first dimension of the velocity field. These are probability density functions of $f(0,z_{0};\theta)$ and $f(0,z_{0})$, where $z_{0}\sim\rho_{0}$. The density functions are further approximated by the histogram of samples. Such techniques for visualization of high dimensional functions have been applied in [33]. We observe that our neural network accurately captures the velocity fields.
@@ 5.2 Flow matching for solving FP equations
Flow matching has emerged as an important problem, which is closely related to generative models [34, 11]. The problem is to simulate the probability density function of a stochastic dynamics:
@@
where $b\colon\mathbb{R}^{+}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a known drift vector field, and $\rho_{0}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is an initial value probability density function. In this example, we also assume the drift function satisfies $b(t,x)=-\nabla V(x)$, which is the negative gradient of some potential function $V(x)\in C_{\text{loc}}^{4}(\mathbb{R}^{d})$. In this case, SDE (33) is the overdamped Langevin dynamic. To simulate the density function of stochastic process $X_{t}$ in (33), we design the following MFC problem. We minimize the objective functional (loss function) as
@@
subject to the FP equation (20). After the score transformation (22), the above second order MFC problem becomes
@@
subject to the transport equation (24). The state dynamic hence becomes $\partial_{t}z_{t}=-\nabla_{z}V(z_{t})-\gamma s_{t}$. This transformation (22) seems to complicate the problem. However, we are able to learn the whole FP equation of the overdamped Langevin dynamic with a given initial distribution $\rho_{0}$ or samples $\{z_{0}^{(n)}\}_{n=1}^{n_{z}}$. More importantly, we can record all intermediate time steps and compute the density function during the entire time domain $[0,t_{\text{end}}]$.
Flow matching for OU process. We first present the flow matching problem for an Ornstein–-Uhlenbeck (OU) process, where the explicit solution is given in A.3. The algorithm is described with details in B.1 and summarized in Algorithm 3. Similar to the previous section, we compare the numerical results for a direct method and the regularized method in Table 4. The errors for the regularized algorithm are significantly smaller.
@@
The numerical results of flow matching for OU process are shown in Figure 3 and 4. Figure 3 shows the density evolution of the process in $1$ dimension, which is similar to the second plot in Figure 1. Our algorithm captures both the change of mean value and the shrink of variance accurately. Figure 4 shows the particle dynamic under the trained probability flow in $2$ dimensions and its comparison with the stochastic OU process. We add level sets of the density function with center $\mu(t)$ and radius $\frac{1}{4}\sigma(t)$, $\frac{1}{2}\sigma(t)$, and $\sigma(t)$ for better visualization, where $\sigma(t):=\det(\Sigma(t))^{\frac{1}{2d}}$ denotes the standard deviation. We pick points that are initialized within the circle with center $\mu(0)$ and radius $\sigma(0)$, and record the trajectories of these particles. According to Figure 4, our algorithm accurately captures the change of mean and variance of the OU process. Additionally, compared with the OU process, our probability flow ODE demonstrates structured behaviors.
Figure 4: Flow matching for OU process in $2$ dimension. First line: particle evolution under the probability flow $z_{t}$ with trained velocity field. Second line: simulation of the stochastic OU process. The circles in the figure are centered at $\mu(t)$, with radius $\frac{1}{4}\sigma(t)$, $\frac{1}{2}\sigma(t)$, and $\sigma(t)$. Our algorithm accurately captures the density evolution of the state dynamic, including the change of mean and variance. The probability flow demonstrates more structured behavior compared with the stochastic OU process.
Double moon example. The example with double moon potential is given by
This example has been computed in [35, 36]. We aim to learn the probability flow ODE of the overdamped Langevin dynamic
@@
We design an algorithm that partitions the time interval into several sub-intervals, which could potentially resolve the issue of long time horizon. The model is trained consecutively over each sub-interval, referred to as the multi-stage splicing method.
Next, we present the detailed implementation of the multi-stage splicing method for our double moon example. In this toy example, we partition the interval into two stages (sub-intervals) $[0,0.2]$ and $[0.2,0.4]$. We apply Algorithm 1 within the first interval. After training, we save the trained network as a warm-start (initialization) for the training in the next stage. We also record the particles $z_{t}^{(n)}$ at terminal time $t=0.2$ in the first stage, which serves as the initial distribution for training in the next stage.
In the second stage $t\in[0.2,0.4]$, the training process is similar. We inherit the network parameter $\theta$ from the previous stage as initialization. Also, the state particles $z_{t}^{(n)}$ at terminal time from the last stage are utilized as initialization distribution for the new stage. Note that starting from the second stage, there is no longer resampling because we only have finite samples at $t=0.2$. As a consequence, we may encounter the problem of overfitting, which is also known as model collapse. In this work, we apply a $L^{2}$ regularization with weight $0.1$ to avoid this issue. If we only have finite samples for the initial distribution, then regularization should also be added in the first stage of training. We summarize this multi-stage splicing method in Algorithm 4.
5.3 An LQ example with an entropy potential cost
We consider an LQ example in this section. This example was also studied in [19]. We minimize the cost functional
$\inf_{v}\int_{0}^{t_{\text{end}}}$ $\int_{\mathbb{R}^{d}}\left(\frac{1}{2}\left|v(t,x)\right|^{2}+\frac{1}{2}|x|^{2}+\beta\log(\rho(t,x))\right)\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t$
$+$ $\int_{\mathbb{R}^{d}}G(x)\rho(t_{\text{end}},x)\,\mathrm{d}x\,,$
subject to
With a score substitution, the problem is equivalent to
$+\int_{\mathbb{R}^{d}}G(x)\rho(t_{\text{end}},x)\,\mathrm{d}x\,,$
subject to
errors $\text{err}_{\rho}$ $\text{err}_{f}$ $\text{err}_{s}$ cost gap
[…]
Table 5: Errors for LQ problem.
The numerical results are also presented in Figure 6. The first, second, and third rows show the results in $1$, $2$, and $10$ dimensions, respectively. The first column is the training curve for the cost functional, which is similar to the first column in Figure 2. Our algorithm gives the correct objective. The second column visualize the density function at terminal time $t_{\text{end}}$ using the data points $\{z^{(n)}_{t_{N_{t}}},\tilde{l}^{(n)}_{t_{N_{t}}}\}_{n=1}^{N_{z}}$. In the $1$ dimension (first row), we plot the density and compare it with true densities directly. In $2$ dimensions, we interpolate the data points and plot the density function $x_{1}\mapsto\rho(t_{\text{end}},(x_{1},0))$ in the figure. In $10$ dimensions, an interpolation is very hard to obtain, so we plot the density in term of the norm of $x$. Our algorithm captures the density function accurately. The third column visualize the score function at terminal time $t_{\text{end}}$ using the data points $\{(z^{(n)}_{t_{N_{t}}},s^{(n)}_{t_{N_{t}}})\}_{n=1}^{N_{z}}$. Again, we plot the score function directly in $1$ dimension. In $2$ and $10$ dimensions, we present a density plot for the first dimension of the score, which gives the probability density function of $\nabla_{z_{1}}\log\rho(t_{\text{end}},z_{t_{\text{end}}})$. We observe that our algorithm captures the true density and score functions accurately.
5.4 Double well potential
In this section we present an MFC example where the terminal cost is a potential function with the double well shape. The formulation is similar to RWPO, with the terminal cost given by
$G(x)=c\left|x-c_{1}\right|^{2}\left|x-c_{2}\right|^{2}\,,$ (35)
where $c\in\mathbb{R}^{+}$ and $c_{1},c_{2}\in\mathbb{R}^{d}$ are the minimizers of $G$. This function $G$ does not depend on the density $\rho$. The initial distribution is still a standard Gaussian distribution, i.e., $z_{0}\sim N(0,I_{d})$. As mentioned before, such problem will demonstrate a bifurcation phenomenon, which is hard to capture. We compute this example in $1$, $2$, and $10$ dimensions, and compute all errors in $1$ and $2$ dimensions. We take $c=\frac{1}{4}$ in (35) and set two centers as an all $-1$ vector for $c_{1}$ and an all $1$ vector for $c_{2}$.
Computing the reference solution. Unlike the example in section 5.1, we do not have an explicit solution. But fortunately, we are able to obtain a reference solution using the kernel formula proposed in [37] (equation (10), (11), and (14)). According to their derivations, the optimal density function is given by
$\rho(t,x)=$ $\left(4\pi\gamma\frac{t(t_{\text{end}}-t)}{t_{\text{end}}}\right)^{-\frac{d}{2}}\cdot$ (36)
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{\exp\left[-\frac{1}{2\gamma}\left(G(z)+\frac{|x-z|^{2}}{2(t_{\text{end}}-t)}+\frac{|x-y|^{2}}{2t}\right)\right]}{\int_{\mathbb{R}^{d}}\exp\left[-\frac{1}{2\gamma}\left(G(\tilde{y})+\frac{|y-\tilde{y}|^{2}}{2t_{\text{end}}}\right)\right]\,\mathrm{d}\tilde{y}}\rho(0,y)\,\mathrm{d}y\,\mathrm{d}z\,.$
Specifically, the density function as terminal time $t_{\text{end}}$ is
In addition, the solution to the classic HJB equation (cf. (81)) is given by
with a terminal condition $\phi(T,x)=-G(x)$. With these expressions, we are able to obtain the score function via $\nabla_{x}\log\rho(t,x)=\nabla_{x}\rho(t,x)/\rho(t,x)$. Then, by Corollary 1 and (82), the optimal velocity is
$f(t,x)=\nabla_{x}\phi(t,x)-\gamma\nabla_{x}\log\rho(t,x)\,.$ (39)
However, when $t\in(0,t_{\text{end}})$, the numerical implementation for $\rho(t,x)$ and $\nabla_{x}\rho(t,x)$ through (36) involves three nested integrations in $\mathbb{R}^{d}$, which could potentially result in large errors. Therefore, we instead consider the errors at $t=0$ and $t_{\text{end}}$, where all expressions are relatively easy. At $t=0$, by (39), we have
At $t=t_{\text{end}}$, we can compute the score function through
Also, the terminal velocity is given by
Next, we apply the Riemann sum to approximate the expressions (40), (41), and (42) to obtain a reference solution in $1$ and $2$ dimensions. First, we define the following function
which has appeared multiple times. We approximated $h(y)$ numerically from the Riemann sum. In $1$ dimension, we use the trapezoid rule with step size $\Delta z=0.01$ to approximate the integrations and truncate the integration within $z\in[-6,6]$ when we compute the reference solution. In $2$ dimensions, similarly, we use a box with size $\Delta z_{1}\times\Delta z_{2}=0.01\times 0.01$ for Riemann sum and truncate the integration within the box $z\in[-6,6]\times[-6,6]$. We compute the value of $h(y)$ on all grid points within $[-4,4]$ in $1$ dimension and $[-4,4]\times[-4,4]$ in $2$ dimensions, and store the values. In this way, integrations w.r.t. $z$ in (40) and (41) are simulated numerically. Next, we further apply Riemann sum for integrations w.r.t. $y$ with the same step size, truncated for $y\in[-4,4]\,\text{or}\,[-4,4]\times[-4,4]$. In this way, we obtain the function for all $x$ on the grid points. The two bounds $6$ and $4$, and the step size $0.01$, are chosen such that the Riemann sums provide reasonable reference solutions for the problem. Finally, we make an interpolation in the dimensions $1$ or $2$, giving us the desired reference solution.
[…]
Figure 7: double well 1d. Left: true density and particle histogram at $t_{\text{end}}$; middle: density evolution compared with true density; right: score function at $t_{\text{end}}$.
[…]
The results are also shown in Figure 8 and 9. Our learned network successfully captures the velocity field $f$ and the score function which is shown in Figure 8. The first and second rows show each dimension of the velocity field at $t=0$ and $t=t_{\text{end}}=1$ and compare them with the true values. The third row shows the score function. The approximated score function is plotted using the data $\{(z^{(n)}_{t_{N_{t}}},s^{(n)}_{t_{N_{t}}})\}_{n=1}^{N_{z}}$, where we made a non-uniform $2d$ interpolation using the build-in function in Scipy. The reference solution is obtained through differentiation of equations (10) (11) and (14) given in [37] and then apply Riemann sum approximation. We observe that our algorithm correctly captures the velocity field and the score function.
The left plot in Figure 9 shows the scattered points at terminal time $t=1$, which concentrate at the two wells of the potential. We also add the level sets of $\rho(1,\cdot)$ at $\{0.01,0.05,0.1,0.2,0.4,0.6\}$ for better visualization. Note that the expression for $\rho(1,\cdot)$ is given by (37) instead of $\frac{1}{Z}\exp(-G(x)/\gamma)$. The plot in the middle shows the particle trajectories under the trained velocity field, which shows a bifurcation phenomenon.
Results in $10$ dimensions. Finally, we also test our algorithm in $10$ dimensions. We do not have a reference solution in this case, but we observe that the particles go into two piles, as is shown in the right plot of Figure 9. In this plot, we project the $10$ dimensional particles into $2$ dimensions for visualization.